OpenStudy (anonymous):

A cylinder carton of oatmeal with radius 3.5 in. is 9 in. tall. If all surfaces except the top are made of cardboard, how much cardboard is used to make the oatmeal carton? Assume no surface overlap. Round your answer to the nearest square inch

OpenStudy (anonymous):

You would do surface area- the top surface area. so surface area is \[\left( R ^{2} \times \pi \right) + \left( diameter \times \pi \times height \right)\] then you would add in the numbers so it would be (3.5^2 *3.14) + (7*3.14*9) = (38.465) + (197.82) = 236.285, if we rounded we would get 237 inches